OpenStudy (anonymous):

3x + 2y = 13 2x + 3y = 12 Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.@.Sam. @april115

OpenStudy (radar):

You could use the elimination method. This method employs a method of changing the coefficient of a variable you choose to eliminate. This can be done by multiplyig and adding. I choose to eliminate the x term. Multiply the first equation thru by -2, getting -6x -4y = -26 Multiply the second equation by 3, getting 6x +9y = 36 Now add the two equations -------------- 5y = 10 or y=2 note the x term just disappeared from the addition. Now substitute the value of y in one of the original equations 3x + 2(2) =13 3x + 4 =13 3x = 13 - 4 = 9 x=3 giving you the final answer (3,2)